# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ 5.1. Convergence of $(\bar{\mu}^{S,N}_{\cdot},\overline{\mathfrak{F}}^{N}(\cdot))$
Remark that for any $i,j\in\mathcal{S}^{N}(0)$, $\kappa^{N}(X_{i}^{N},X_{j}^{N})\cdot\gamma^{N}(X_{i}^{N},X_{j}^{N})=\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})/N\leq\omega^{*}/N$. Recalling the definitions given in Assumption 3.5, we thus deduce
@@
Recalling (5.31) and (5.32), this entails ¯$\overline{}\mathfrak{U}^{N}\leq\overline{}\mathfrak{V}^{N}$ a.s. and concludes the proof of Lemma 5.6. ∎
For the upper-bound of ¯$\overline{}\mathfrak{V}_{i}^{N,1}$ and ¯$\overline{}\mathfrak{V}_{i}^{N,0}$ as defined in respectively (5.9) and (5.13), the proof of the next lemma follows similar principles as in the previous one.
@@
The following upper-bound holds for any $t>0$ with the sequence (¯$(\bar{}\mathfrak{V}^{N})_{N\geq 1}$ defined in (5.27):
@@
We treat this component similarly as $\mathfrak{U}^{N}$, first thanks to the Cauchy-Schwartz inequality:
$\begin{split}\mathbb{E}_{0}^{N}\Big{[}\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\big{|}\overline{}\mathfrak{V}_{i}^{N,1}(t)\big{|}\Big{]}&\leq\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\sqrt{\mathbb{E}_{0}^{N}\Big{[}\big{(}\overline{}\mathfrak{V}_{i}^{N,1}(t)\big{)}^{2}\Big{]}}\\ &\leq\sqrt{\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\mathbb{E}_{0}^{N}\Big{[}\big{(}\overline{}\mathfrak{V}_{i}^{N,1}(t)\big{)}^{2}\Big{]}},\end{split}$
Recalling (5.9), we note that the random variables ¯$\overline{}\mathfrak{V}_{i}^{N,1}(t)$ are centered conditionally on $\mathcal{F}^{N}_{0}$, so that the term under the square root is actually a variance. ¯$\overline{}\mathfrak{V}_{i}^{N,1}(t)$ is a sum of r.v.’s which are orthogonal in $L^{2}(\Omega,\mathbb{P}_{0}^{N})$. Thus, for any $i\in\mathcal{S}^{N}(0)$,
@@
Concerning the sequence (¯$(\overline{}\mathfrak{V}_{i}^{N,0}(t))$, we first recall the following identity for any $j\in\mathcal{I}^{N}(0)$ as part of Assumption 2.6:
@@
Since $\omega^{N}(i,j)$ and $\lambda_{j}^{N}(A_{j}^{N}(0)+t)$ are independent conditionally on $\mathcal{F}^{N}_{0}$, we deduce that ¯$\overline{}\mathfrak{V}_{i}^{N,0}(t)$ is also conditionally centered, whatever $i\in\mathcal{S}^{N}(0)$ and $t\geq 0$. We can then exploit the same argument for ¯$\overline{}\mathfrak{V}_{i}^{N,0}(t)$ as for ¯$\overline{}\mathfrak{V}^{N,1}_{i}(t)$, thanks to the Cauchy-Schwartz inequality and replacing (5.35) by
@@
Lemma 5.7 is therefore concluded with ¯$\overline{}\mathfrak{V}^{N}$ in (5.27). ∎
¯$\overline{}\mathfrak{L}^{N,1}$ and ¯$\overline{}\mathfrak{L}^{N,0}$ as defined respectively in (5.10) and (5.14) are quite directly upper-bounded.
@@
By virtue of Assumptions 2.6 and 3.1:
|¯$\Big{|}\overline{}\mathfrak{L}^{N,1}(t,x)\Big{|}+\Big{|}\overline{}\mathfrak{L}^{N,0}(t,x)\Big{|}\leq\lambda^{*}\cdot\|\bar{\omega}^{N}-\bar{\omega}\|_{\infty}\,$
@@
In order to finally deal with the upper-bound of both ¯$\overline{}\mathfrak{E}^{N,1}$ and ¯$\overline{}\mathfrak{E}^{N,0}$, defined respectively in (5.12) and (5.15), we exploit the following proposition which will be proved in Appendix A.
@@
We are now ready for the estimate of ¯$\overline{}\mathfrak{E}^{N,1}$ and ¯$\overline{}\mathfrak{E}^{N,0}$ provided in the next lemma.
@@
¯$\overline{}\mathfrak{E}^{N}(t):=\langle\bar{\mu}^{S,N}_{0},|\overline{}\mathfrak{E}^{N,1}(t,.)|+|\overline{}\mathfrak{E}^{N,0}(t,.)|\rangle\,.$
@@
$\begin{split}\mu^{N}:=\bar{\mu}^{S,N}_{0},\quad\mu:=\bar{\mu}^{S}_{0},\qquad&k(x,x^{\prime},a^{\prime})=\begin{cases}&\bar{\omega}(x,x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F_{x^{\prime}}^{c}(a^{\prime})}\quad\text{ if }F_{x^{\prime}}^{c}(a^{\prime}+t)>0\\ &0\hskip 99.58464pt\text{otherwise,}\end{cases}\\ \nu^{N}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime}):=\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime}),\quad&\nu(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime}):=\bar{\mu}^{I}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\,.\end{split}$
Assumption 3.1 states that $\mu^{N}$ and $\nu^{N}$ converge weakly to respectively $\mu$ and $\nu$. Exploiting first the fact that $a\mapsto F^{c}_{x^{\prime}}(a)$ is non increasing, and next Assumption 2.6, we have that for all $x^{\prime}\in\mathbf{\mathbb{X}},a^{\prime},t\geq 0$,
$\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\leq\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime}+t)}\leq\lambda^{\ast}\,.$
We recall also the uniform upper-bound $\omega^{*}$ of $\bar{\omega}^{N}$, as defined just after Lemma 3.4. $\omega^{*}\cdot\lambda^{*}$ thus defines a global upper-bound of $k$.
Note that $\bar{\lambda}(x^{\prime},a^{\prime})<\lambda^{*}\cdot F_{x^{\prime}}^{c}(a^{\prime})$ as long as $F_{x^{\prime}}^{c}(a^{\prime})>0$, by virtue of Assumption 2.6. By virtue of Assumptions 2.4, 2.6 and 3.3, $k$ is $(\mu\otimes\nu)$ a.e. continuous, those points $(x,x^{\prime},a^{\prime})$ that satisfy $F^{c}_{x^{\prime}}(a^{\prime})=0$ being handled specifically thanks to the above property. Indeed, thanks to the second part of Assumption 3.3 for $(\mu\otimes\nu)$ almost every such realization, since $t>0$, we know that for any $(\check{x},\check{x}^{\prime},\check{a}^{\prime})$ in a small enough neighborhood of $(x,x^{\prime},a^{\prime})$ that $F^{c}_{\check{x}^{\prime}}(\check{a}^{\prime}+t)=0$, thus $\bar{\lambda}(\check{x}^{\prime},\check{a}^{\prime}+t)=0$ and $k(\check{x},\check{x}^{\prime},\check{a}^{\prime})=0$.
@@
$\mu^{N}=\nu^{N}:=\bar{\mu}^{S,N}_{0},\quad\mu=\nu=\bar{\mu}^{S}_{0},\quad k(x,x^{\prime})=\bar{\omega}(x,x^{\prime})\cdot\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\,.$
The fact that $\mu^{N}$ and $\nu^{N}$ converge weakly in probability to $\mu=\nu$ follows from Assumption 3.1. $k$ is bounded under Assumptions 2.2 and 3.3. To check that $k$ is $(\mu\otimes\nu)$ a.e. continuous, we exploit the following alternative expression:
$k(x,x^{\prime})=\bar{\omega}(x,x^{\prime})\int_{0}^{t}\bar{\lambda}(x^{\prime},t-s)\cdot\overline{\mathfrak{F}}(s,x^{\prime})\cdot\exp\Big{[}-\int_{0}^{s}\overline{\mathfrak{F}}(r,x^{\prime})\mathrm{d}r\Big{]}\,,$
derived with the same argument as for (5.24). Thanks to Assumptions 2.2 and 3.3 and Lemma 4.2, we then conclude that $k$ is $(\mu\otimes\nu)$ a.e. continuous. ∎
@@
As $N\to\infty$, ¯$\overline{}\mathfrak{D}^{N}(t)$ defined in (5.6) converges in probability to 0 locally uniformly in $t$.
@@
First note that it suffices to prove the convergence for any fixed $t$. The locally uniform convergence then follows from Lemma A.1 in Appendix A since $:t\mapsto\overline{}\mathfrak{D}^{N}(t)$ is a.s. non-decreasing for any $N$. For any $i$ and $t$,
@@
¯$\overline{}\mathfrak{D}^{N}(t)\leq\mathbb{E}_{0}^{N}\Big{[}\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\int_{0}^{t}|\overline{\mathfrak{F}}_{i}^{N}(s)-\overline{\mathfrak{F}}(s,X_{i}^{N})|\mathrm{d}s\Big{]}\,.$ (5.38)
Starting from (5.38), for any $i$ and $t$, we decompose the integrand into seven terms according to Lemma 5.4. Five among these terms are first treated thanks to Lemmas 5.6, 5.7 and 5.8, so that there exists $C>0$, ¯$\overline{}\mathfrak{V}^{N}=\lambda^{*}\cdot\sqrt{\Upsilon^{N}+\omega^{*}\cdot\bar{\gamma}^{N}}$ and ¯$\overline{}\mathfrak{L}^{N}=\lambda^{*}\cdot\|\bar{\omega}^{N}-\bar{\omega}\|_{\infty}$, the later two converging in probability to zero by virtue respectively of Assumptions 3.5 and 3.3, such that
¯$\overline{}\mathfrak{D}^{N}(t)\leq C\cdot\int_{0}^{t}\overline{}\mathfrak{D}^{N}(s)+t\cdot(2\overline{}\mathfrak{V}^{N}+\overline{}\mathfrak{L}^{N})+\int_{0}^{t}\langle\bar{\mu}^{S,N}_{0},|\overline{}\mathfrak{E}^{N,1}(s,.)|+|\overline{}\mathfrak{E}^{N,0}(s,.)|\rangle\mathrm{d}s\,.$ (5.39)
Thanks to Lemma 5.10, the integrand $(\langle\bar{\mu}^{S,N}_{0},|\overline{}\mathfrak{E}^{N,1}(s,.)|+|\overline{}\mathfrak{E}^{N,0}(s,.)|\rangle)_{s\in[0,t]}$ converges to 0 pointwise in probability for any $s>0$. By virtue of Assumption 2.6 and Lemma 3.4, for the defining properties of $\lambda^{*}$ and $\omega^{*}$, we obtain that
|¯$|\overline{}\mathfrak{E}^{N,1}(s,x)|\vee|\overline{}\mathfrak{E}^{N,0}(s,.)|\leq\omega^{*}\,\lambda^{*},$
@@
With (5.39), we are therefore in situation to apply Gronwall’s inequality and conclude Proposition 5.11 in that ¯$\overline{}\mathfrak{D}^{N}(t)$ converges to 0 in probability for any $t$ as $N\to\infty$. ∎
Before proceeding, let us state a result in Proposition 5.12 which is exactly Theorem II.4.1 in [48] and will be needed in the next proof. Let us generally consider a Polish space $\mathcal{X}$ and the set $\mathcal{C}_{b}(\mathcal{X})$ of bounded continuous functions on $\mathcal{X}$. We say that a subset $\mathfrak{M}$ of $\mathcal{C}_{b}(\mathcal{X})$ is separating if $(i)$ it includes the constant function $:x\mapsto 1$ and $(ii)$ it discriminates elements of $\mathcal{M}_{1}(\mathcal{X})$, in that for any $(\nu,\nu^{\prime})\in\mathcal{M}_{1}(\mathcal{X})^{2}$, $\nu=\nu^{\prime}$ is equivalent to the property that $\langle\nu,\phi\rangle=\langle\nu,\phi\rangle$ for any $\phi\in\mathfrak{M}$. It is classical that $\mathcal{C}_{b}(\mathcal{X})$ itself is separating.
@@
Concerning Point $(b)$, let us consider any $\varphi\in C_{b}(\mathbf{\mathbb{X}})$, $\varphi\geq 0$. Recall that from Lemma 5.1 $\langle\widetilde{\mu}^{(S,N)}_{t},\varphi\rangle$ converges in probability to $\langle\bar{\mu}^{S}_{t},\varphi\rangle$ for any $t>0$. Moreover, by definition of $\widetilde{\mu}^{(S,N)}$ and ¯$\overline{}\mathfrak{D}^{N}$ in respectively (5.2) and (5.6):
@@
The above arguments directly entail the following pointwise in time convergence result on the potential force of infection.
@@
The following convergence to 0 in conditional expectation holds for any $t>0$:
@@
where $C,\lambda^{*}<\infty$ while ¯$\overline{}\mathfrak{D}^{N}(t)$, ¯$\overline{}\mathfrak{V}^{N}$, $\|\bar{\omega}^{N}-\bar{\omega}\|_{\infty}$ and ¯$\overline{}\mathfrak{E}^{N}(t)$ all tend to 0 thanks respectively to Proposition 5.11, Assumption 3.5, Assumption 3.3, and Lemma 5.10. This concludes the proof of Proposition 5.14. ∎
@@
[…]
The test function $\varphi$ evaluates here the convergence of a distribution on $\mathbf{\mathbb{X}}$ that we call the activated force of infection. The measure on the left-hand side can be interpreted as minus the derivative of the process $\bar{\mu}^{S}$ at time $t$.

### v2

@@ 5.1. Convergence of $(\bar{\mu}^{S,N}_{\cdot},\overline{\mathfrak{F}}^{N}(\cdot))$
Remark that from (2.6), for any $i,j\in\mathcal{S}^{N}(0)$, $\kappa^{N}(X_{i}^{N},X_{j}^{N})\cdot\gamma^{N}(X_{i}^{N},X_{j}^{N})=\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})/N\leq\omega^{*}/N$. Recalling the definitions given in Assumption 3.4, we thus deduce
@@
Recalling (5.31) and (5.32), this entails $\overline{\mathfrak{U}}^{N}\leq\overline{\mathfrak{V}}^{N}$ a.s. and concludes the proof of Lemma 5.6. ∎
For the upper-bound of $\overline{\mathfrak{V}}_{i}^{N,1}$ and $\overline{\mathfrak{V}}_{i}^{N,0}$ as defined in respectively (5.9) and (5.13), the proof of the next lemma follows similar principles as in the previous one.
@@
The following upper-bound holds for any $t>0$ with the sequence $(\bar{}\mathfrak{V}^{N})_{N\geq 1}$ defined in (5.27):
@@
We treat this component in the same way as $\mathfrak{U}^{N}$, starting with the Cauchy-Schwartz inequality:
$\begin{split}\mathbb{E}_{0}^{N}\Big[\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\big|\overline{\mathfrak{V}}_{i}^{N,1}(t)\big|\Big]\leq\sqrt{\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\mathbb{E}_{0}^{N}\Big[\big(\overline{\mathfrak{V}}_{i}^{N,1}(t)\big)^{2}\Big]},\end{split}$
Recalling (5.9), we note that the random variables $\overline{\mathfrak{V}}_{i}^{N,1}(t)$ are centered conditionally on $\mathcal{F}^{N}_{0}$, so that the term under the square root is actually a variance. $\overline{\mathfrak{V}}_{i}^{N,1}(t)$ is a sum of r.v.’s which are orthogonal in $L^{2}(\Omega,\mathbb{P}_{0}^{N})$. Thus, for any $i\in\mathcal{S}^{N}(0)$,
@@
Concerning the sequence $(\overline{\mathfrak{V}}_{i}^{N,0}(t))$, we first recall the following identity for any $j\in\mathcal{I}^{N}(0)$ as part of Assumption 2.6:
@@
Since $\omega^{N}(i,j)$ and $\lambda_{j}^{N}(A_{j}^{N}(0)+t)$ are independent conditionally on $\mathcal{F}^{N}_{0}$, we deduce that $\overline{\mathfrak{V}}_{i}^{N,0}(t)$ is also conditionally centered, whatever $i\in\mathcal{S}^{N}(0)$ and $t\geq 0$. We can then exploit the same argument for $\overline{\mathfrak{V}}_{i}^{N,0}(t)$ as for $\overline{\mathfrak{V}}^{N,1}_{i}(t)$, thanks to the Cauchy-Schwartz inequality and replacing (5.35) by
@@
Lemma 5.7 is therefore concluded with $\overline{\mathfrak{V}}^{N}$ in (5.27). ∎
$\overline{\mathfrak{L}}^{N,1}$ and $\overline{\mathfrak{L}}^{N,0}$ as defined respectively in (5.10) and (5.14) are quite directly upper-bounded.
@@
By virtue of Assumptions 2.4, 2.6 and 3.1:
$\Big|\overline{\mathfrak{L}}^{N,1}(t,x)\Big|+\Big|\overline{\mathfrak{L}}^{N,0}(t,x)\Big|\leq\lambda^{*}\cdot\|\bar{\omega}^{N}-\bar{\omega}\|_{\infty}\,$
@@
In order to finally deal with the upper-bound of both $\overline{\mathfrak{E}}^{N,1}$ and $\overline{\mathfrak{E}}^{N,0}$, defined respectively in (5.12) and (5.15), we exploit the following proposition which will be proved in Appendix A.
@@
We are now ready for the estimate of $\overline{\mathfrak{E}}^{N,1}$ and $\overline{\mathfrak{E}}^{N,0}$ provided in the next lemma.
@@
$\overline{\mathfrak{E}}^{N}(t):=\langle\bar{\mu}^{S,N}_{0},|\overline{\mathfrak{E}}^{N,1}(t,.)|+|\overline{\mathfrak{E}}^{N,0}(t,.)|\rangle\,.$
@@
$\begin{split}\mu^{N}:=\bar{\mu}^{S,N}_{0},\quad\mu:=\bar{\mu}^{S}_{0},\qquad&k_{t}^{0}(x,x^{\prime},a^{\prime})=\begin{cases}&\bar{\omega}(x,x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F_{x^{\prime}}^{c}(a^{\prime})}\quad\text{ if }{\color[rgb]{0,0.3,0}F_{x^{\prime}}^{c}(a^{\prime})>0,}\\ &0\hskip 99.58464pt\text{otherwise,}\end{cases}\\ \nu^{N,0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime}):=\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime}),\quad&\nu^{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime}):=\bar{\mu}^{I}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\,.\end{split}$
Assumption 3.1 states that $\mu^{N}$ and $\nu^{N,0}$ converge weakly to respectively $\mu$ and $\nu^{0}$. We recall the upper-bound of $\bar{\lambda}(x^{\prime},a^{\prime}+t)/F^{c}_{x^{\prime}}(a^{\prime})$ by $\lambda^{*}$ as a consequence of Assumptions 2.4 and 2.6. We recall also the uniform upper-bound $\omega^{*}$ of $\bar{\omega}^{N}$, as defined in Assumption 3.3. $\omega^{*}\cdot\lambda^{*}$ is thus a global upper-bound of $k_{t}^{0}$.
By virtue of Assumption 3.2 and since $t>0$, the following property holds for $(\mu\otimes\nu^{0})$ almost every triplet $(x,x^{\prime},a^{\prime})$ such that $F^{c}_{x^{\prime}}(a^{\prime})=0$: $F^{c}_{\check{x}^{\prime}}(\check{a}^{\prime}+t)=0$ holds for any $(\check{x},\check{x}^{\prime},\check{a}^{\prime})$ in a small enough neighborhood of $(x,x^{\prime},a^{\prime})$, thus also $\bar{\lambda}(\check{x}^{\prime},\check{a}^{\prime}+t)=0$ and $k(\check{x},\check{x}^{\prime},\check{a}^{\prime})=0$.
The following property holds on the other hand, for $(\mu\otimes\nu^{0})$ almost every triplet $(x,x^{\prime},a^{\prime})$ such that $F^{c}_{x^{\prime}}(a^{\prime})>0$: $F^{c}_{\check{x}^{\prime}}(\check{a}^{\prime})>0$ holds for any $(\check{x},\check{x}^{\prime},\check{a}^{\prime})$ in a small enough neighborhood of $(x,x^{\prime},a^{\prime})$. Still by virtue of Assumption 3.2 on $\bar{\lambda}$ and $\bar{\mu}_{0}^{I}$, the function $(x^{\prime},a^{\prime})\mapsto\bar{\lambda}(x^{\prime},a^{\prime}+t)$ is $\nu^{0}$-a.e. continuous. We thus conclude that $k_{t}^{0}$ is $(\mu\otimes\nu^{0})$-a.e. continuous.
Proposition 5.9 does therefore entail that $\langle\bar{\mu}^{S,N}_{0},|\overline{\mathfrak{E}}^{N,0}(t,.)|\rangle$ tends to zero.
@@
$\mu^{N}=\nu^{N,1}:=\bar{\mu}^{S,N}_{0},\quad\mu=\nu^{1}=\bar{\mu}^{S}_{0},\quad k^{1}_{t}(x,x^{\prime})=\bar{\omega}(x,x^{\prime})\cdot\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\,.$
The fact that $\mu^{N}$ and $\nu^{N,1}$ converge weakly in probability to $\mu=\nu^{1}$ follows from Assumption 3.1. $k_{t}^{1}$ is bounded under Assumptions 2.2 and 3.2. To check that $k^{1}_{t}$ is $(\mu\otimes\nu^{1})$ a.e. continuous, we exploit the following alternative expression:
$k_{t}^{1}(x,x^{\prime})=\bar{\omega}(x,x^{\prime})\int_{0}^{t}\bar{\lambda}(x^{\prime},t-s)\cdot\overline{\mathfrak{F}}(s,x^{\prime})\cdot\exp\Big[-\int_{0}^{s}\overline{\mathfrak{F}}(r,x^{\prime})\mathrm{d}r\Big]\,\mathrm{d}s\,$,
derived with the same argument as for (5.24). Thanks to Assumptions 2.2 and 3.2 and Lemma 4.2, we can conclude that $k^{1}$ is $(\mu\otimes\nu^{1})$ a.e. continuous, as stated in Lemma A.3 in Appendix A.
Proposition 5.9 does therefore entail that $\langle\bar{\mu}^{S,N}_{0},|\overline{\mathfrak{E}}^{N,1}(t,.)|\rangle$ tends to zero. ∎
@@
As $N\to\infty$, $\overline{\mathfrak{D}}^{N}(t)$ defined in (5.6) converges in probability to 0 locally uniformly in $t$.
@@
First note that it suffices to prove the convergence for any fixed $t$. The locally uniform convergence then follows from Lemma A.1 in Appendix A since $t\mapsto\overline{\mathfrak{D}}^{N}(t)$ is a.s. non-decreasing for any $N$. For any $i$ and $t$,
@@
$\overline{\mathfrak{D}}^{N}(t)\leq\mathbb{E}_{0}^{N}\Big[\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\int_{0}^{t}|\overline{\mathfrak{F}}_{i}^{N}(s)-\overline{\mathfrak{F}}(s,X_{i}^{N})|\mathrm{d}s\Big]\,.$ (5.38)
Starting from (5.38), for any $i$ and $t$, we decompose the integrand into seven terms according to Lemma 5.4. Five among these terms are first treated thanks to Lemmas 5.6, 5.7 and 5.8, so that there exists $C>0$, $\overline{\mathfrak{V}}^{N}=\lambda^{*}\cdot\sqrt{\Upsilon^{N}+\omega^{*}\cdot\bar{\gamma}^{N}}$ and $\overline{\mathfrak{L}}^{N}=\lambda^{*}\cdot\|\bar{\omega}^{N}-\bar{\omega}\|_{\infty}$, the later two converging in probability to zero by virtue respectively of Assumptions 3.4 and 3.2, such that
$\overline{\mathfrak{D}}^{N}(t)\leq C\cdot\int_{0}^{t}\overline{\mathfrak{D}}^{N}(s)+t\cdot(2\overline{\mathfrak{V}}^{N}+\overline{\mathfrak{L}}^{N})+\int_{0}^{t}\langle\bar{\mu}^{S,N}_{0},|\overline{\mathfrak{E}}^{N,1}(s,.)|+|\overline{\mathfrak{E}}^{N,0}(s,.)|\rangle\mathrm{d}s\,.$ (5.39)
Thanks to Lemma 5.10, the integrand $(\langle\bar{\mu}^{S,N}_{0},|\overline{\mathfrak{E}}^{N,1}(s,.)|+|\overline{\mathfrak{E}}^{N,0}(s,.)|\rangle)_{s\in[0,t]}$ converges to 0 pointwise in probability for any $s>0$. By virtue of Assumptions 2.6 and 3.3, for the defining properties of $\lambda^{*}$ and $\omega^{*}$, we obtain that
$|\overline{\mathfrak{E}}^{N,1}(s,x)|\vee|\overline{\mathfrak{E}}^{N,0}(s,.)|\leq\omega^{*}\,\lambda^{*},$
@@
With (5.39), we are therefore in situation to apply Gronwall’s inequality and conclude Proposition 5.11 in that $\overline{\mathfrak{D}}^{N}(t)$ converges to 0 in probability for any $t$ as $N\to\infty$. ∎
Before proceeding, let us state a result which is exactly Theorem II.4.1 in [45] and will be needed in the next proof. Let us generally consider a Polish space $\mathcal{X}$. We say that a subset $\mathfrak{M}$ of $\mathcal{C}_{b}(\mathcal{X})$ is separating if $(i)$ it includes the constant function $x\mapsto 1$ and $(ii)$ it discriminates elements of $\mathcal{M}_{1}(\mathcal{X})$, in that for any $(\nu,\nu^{\prime})\in\mathcal{M}_{1}(\mathcal{X})^{2}$, $\nu=\nu^{\prime}$ is equivalent to the property that $\langle\nu,\phi\rangle=\langle\nu,\phi\rangle$ for any $\phi\in\mathfrak{M}$. It is well known that $\mathcal{C}_{b}(\mathcal{X})$ itself is separating. A sequence of random elements of $\mathcal{D}(\mathbb{R}_{+},\mathcal{X})$ is said to be C-tight if it is tight and any limit of a converging subsequence is a.s. continuous.
@@
Concerning Point $(b)$, let us consider any $\varphi\in\mathcal{C}_{b,+}(\mathbf{\mathbb{X}})$. Recall that from Lemma 5.1 $\langle\widetilde{\mu}^{(S,N)}_{t},\varphi\rangle$ converges in probability to $\langle\bar{\mu}^{S}_{t},\varphi\rangle$ for any $t>0$. Moreover, by definition of $\widetilde{\mu}^{(S,N)}$ and $\overline{\mathfrak{D}}^{N}$ in respectively (5.2) and (5.6):
@@
The above arguments directly entail the following pointwise in time convergence result on the force of infection.
@@
The following convergence to 0 holds a.s. for any $t>0$:
@@
where $C,\lambda^{*}<\infty$ while $\overline{\mathfrak{D}}^{N}(t)$, $\overline{\mathfrak{V}}^{N}$, $\|\bar{\omega}^{N}-\bar{\omega}\|_{\infty}$ and $\overline{\mathfrak{E}}^{N}(t)$ all tend to 0 thanks respectively to Proposition 5.11, Assumption 3.4, Assumption 3.2, and Lemma 5.10. This concludes the proof of Proposition 5.14. ∎
@@
$\sum_{i\leq N}\overline{\mathfrak{F}}_{i}^{N}(t)\,\mathbbm{1}_{\left\{D_{i}^{N}(t)=0\right\}}\,\varphi(X^{N}_{i})\to\langle\overline{\mathfrak{F}}(t)\bar{\mu}^{S}_{t},\varphi\rangle\,.$
The test function $\varphi$ evaluates here the convergence of a distribution on $\mathbf{\mathbb{X}}$ that we call the activated force of infection. The measure on the left-hand side can be interpreted as minus the derivative of the process $\bar{\mu}^{N,S}$ at time $t$.